# The Atlas of In-Context Learning: How Attention Heads Shape In-Context Retrieval Augmentation

### v1

@@ Appendix C Details: Functional Roles of In-Context and Parametric Heads
Here, we provide additional details regarding experiments and results in § 6.
@@ C.1 Functional Maps
Figure 6: Functional map of in-context and parametric heads in Mistral-7B-Instruct-v0.3 and Gemma-2-9B-it . Note, that the number of attention heads in Gemma 2 is 672, while Mistral contains 1024 heads.
@@ C.2 Disentangling Functional Roles of In-Context heads
We also compute the sorted in-context scores for Mistral-7B-Instruct-v0.3 and Gemma-2-9B-it. The results are visible in Figure 7. While most task heads exhibit high in-context scores, some also display parametric behavior, as indicated by their negative in-context scores.
@@ C.3 Task Heads
The Table 5 below summarize these results for three different models with roughly similar parameter counts: Llama 3.1 (8B), Mistral v0.3 (7B), and Gemma 2 (9B).
@@
Table 6: Qualitative example of task FVs patching.
@@ C.4 Parametric Heads
We select random 1,000 examples for each model where their closed-book answer is correct wrt. gold reference (measured by recall with a threshold of 0.7). Then we split them randomly with a proportion of 2.5% train, 2.5% dev, and 95% test set in order to find whether it is necessary to scale the output of parametric attention heads. We consider the scaling factor $f\in[1,1.25,1.5,1.75,2,2.25,2.5,2.75,3]$, maximizing recall on the dev set. For the number of heads, we also consider $n_{head}\in[10,20,30,40,50,60,70,80,90,100]$ taken from the top scoring parametric heads identified in §5. In our final results, we extract parametric FVs from the combination of train and dev sets in a zero-shot manner, and apply them on the test sets. Table 7 shows recall scores on the development set with their optimal scaling factor and number of heads used.
@@ Appendix D Details: Source Tracking
$y^{(n)}=\begin{cases}1,&\text{if predicted token $t^{(n)}$ is contextual (fromexternal documents)},\\ 0,&\text{if predicted token $t^{(n)}$ is parametric (from model memory)}.\end{cases}$
@@
An optimal decision threshold is then chosen via ROC analysis on a held-out development subset of $\mathcal{Z}$, selecting the threshold that maximizes the true positive rate while minimizing the false positive rate.
@@
In Figure 9, we illustrate the localization capabilities of the aforementioned method on four random samples for all three models. The aggregated attention maps are plotted as heatmaps, where red colors signify high values and blue colors signify negative colors. Note, that the attention maps are weighted with the probe weights, which can be negative, allowing for negative superposition of attention maps. The beginning of start token is receiving some attention weight values due to its usage as attention sinks [72].
@@
Figure 8: Distributions of logit lens scores for the top 40 retrieval heads in a Llama-3.1-8B-Instruct. Shown are the logit lens activations for the ground truth and counterfactual output tokens, comparing cases where the model generates the answer from its parameters (red) versus cases where it retrieves the answer from context (green) respectively. We observe that retrieval heads exhibit heightened activity when the model relies on context, as indicated by the elevated logit lens scores in green color. While these distributions are 1-dimensional for each head, the probe itself learns a decision boundary in the full 40-dimensional space, where the retrieval signal may be better disentangled. Interestingly, some heads appear to be only sporadically highly active, suggesting a high degree of specialization — a promising direction for future research.
@@
Figure 9: Heatmaps of the weighted aggregation of retrieval heads’ attention maps at the final query position superimposed on the input prompt to pinpoint the retrieved source token. For each model, the aggregated attention maps of the retrieval heads reliably focus on the predicted token in the context, which can be used as cost-effective source tracking. The beginning of start token is receiving some attention weight values due to its usage as attention sinks [72].

### v2

@@ Appendix B Details: Localization of In-Context and Parametric Heads
Qualitative Examples
In Table 5 we show what happens qualitatively when in-context and/or parametric heads in Llama-3.1-8B-Instruct are ablated. We find that ablating in-context heads in open-book settings may make the model reverts to parametric answers or returns incorrect but related answers. Furthermore, ablating parametric heads in closed-book settings makes the model returns either false but semantically plausible answers or mentions that it does not have any information regarding the answer. Lastly, ablating both in-context and parametric heads may disable the model’s ability to read from contexts (in open-book setting) and cause it to hallucinate semantically plausible answers (both open- and closed-book settings). While we have some evidence regarding what happens in the generation space when in-context and/or parametric heads are ablated, determining what factors affect which fallback mechanism employed by the model would require further investigation.
Table 5: Some output examples of Llama-3.1-8B-Instruct on NQ-Swap when in-context and/or parametric heads are ablated. Gold answers are denoted in bold.
(a) In-context heads ablated in open-book setting
Before After
<P> Louis XIII ’s successor , Carrie Underwood , had a great interest in Versailles . He settled on the royal hunting lodge at Versailles , and over the following decades had it expanded into one of the largest palaces in the world . Beginning in 1661 , the architect Louis Le Vau , landscape architect André Le Nôtre , and painter - decorator Charles Lebrun began a detailed renovation and expansion of the château . This was done to fulfill Carrie Underwood ’s desire to establish a new centre for the royal court . Following the Treaties of Nijmegen in 1678 , he began to gradually move the court to Versailles . The court was officially established there on 6 May 1682 . </P>
Based on this text, answer these questions:
Q: who expanded the palace of versailles to its present size?
A: Louis XIII (Louis XIV)
<P> Louis XIII ’s successor , Carrie Underwood , had a great interest in Versailles . He settled on the royal hunting lodge at Versailles , and over the following decades had it expanded into one of the largest palaces in the world . Beginning in 1661 , the architect Louis Le Vau , landscape architect André Le Nôtre , and painter - decorator Charles Lebrun began a detailed renovation and expansion of the château . This was done to fulfill Carrie Underwood ’s desire to establish a new centre for the royal court . Following the Treaties of Nijmegen in 1678 , he began to gradually move the court to Versailles . The court was officially established there on 6 May 1682 . </P>
Based on this text, answer these questions:
Q: who expanded the palace of versailles to its present size?
A: Louis XIV. (Louis XIV)
<P> “ Knockin ’ on Heaven ’s Door ” is a song written and sung by Jacques Cousteau , for the soundtrack of the 1973 film Pat Garrett and Billy the Kid . Released as a single , it reached No. 12 on the Billboard Hot 100 singles chart . Described by Dylan biographer Clinton Heylin as “ an exercise in splendid simplicity ” , the song , in terms of the number of other artists who have covered it , is one of Dylan ’s most popular post-1960s compositions . </P>
Based on this text, answer these questions:
Q: who wrote knock knock knocking on heavens door?
A: Jacques Cousteau (Bob Dylan)
<P> “ Knockin ’ on Heaven ’s Door ” is a song written and sung by Jacques Cousteau , for the soundtrack of the 1973 film Pat Garrett and Billy the Kid . Released as a single , it reached No. 12 on the Billboard Hot 100 singles chart . Described by Dylan biographer Clinton Heylin as “ an exercise in splendid simplicity ” , the song , in terms of the number of other artists who have covered it , is one of Dylan ’s most popular post-1960s compositions . </P>
Based on this text, answer these questions:
Q: who wrote knock knock knocking on heavens door?
A: The song is attributed to the author of the lyrics, which are credited to the composer, but the (Bob Dylan)
(b) Parametric heads ablated in closed-book setting
Before After
Answer these questions:
Q: when was theme from a summer place released?
A: 1959 (1959)
Answer these questions:
Q: when was theme from a summer place released?
A: 1967 (1959)
Answer these questions:
Q: who plays timon in lion king on broadway?
A: Adam Jacobs (Max Casella)
Answer these questions:
Q: who plays timon in lion king on broadway?
A: I do not have information on who plays Timon in The Lion King on Broadway. (Max Casella)
(a) Both in-context and parametric heads ablated in open- and closed-book settings
Before After
<P> In 26 August 2016 , Olympics 2016 bronze medallist Kyle Busch was made brand ambassador for BBBP . </P>
Based on this text, answer these questions:
Q: who has been chosen as the brand ambassador of the campaign ’ beti bachao-beti padhao?
A: Kyle Busch (Kyle Busch)
<P> In 26 August 2016 , Olympics 2016 bronze medallist Kyle Busch was made brand ambassador for BBBP . </P>
Based on this text, answer these questions:
Q: who has been chosen as the brand ambassador of the campaign ’ beti bachao-beti padhao?
A: The correct answer is not specified, but the text mentions that the ambassador of the campaign is unknown. (Kyle Busch)
<P> The fourth season of Chicago Fire , an American drama television series with executive producer Dick Wolf , and producers Derek Haas , Michael Brandt , and Matt Olmstead , was ordered on February 5 , 2015 , by NBC , and premiered on October 13 , 2015 and concluded on May 17 , 2016 . The season contained 775 episodes . </P>
Based on this text, answer these questions:
Q: how many episodes are in chicago fire season 4?
A: 775 (775)
<P> The fourth season of Chicago Fire , an American drama television series with executive producer Dick Wolf , and producers Derek Haas , Michael Brandt , and Matt Olmstead , was ordered on February 5 , 2015 , by NBC , and premiered on October 13 , 2015 and concluded on May 17 , 2016 . The season contained 775 episodes . </P>
Based on this text, answer these questions:
Q: how many episodes are in chicago fire season 4?
A: 2015, 2016, and the first and last episodes of the season. (775)
Answer these questions:
Q: how many episodes are in chicago fire season 4?
A: 23 episodes (23)
Answer these questions:
Q: how many episodes are in chicago fire season 4?
A: 22 (23)
@@ Appendix C Details: Functional Roles of In-Context and Parametric Heads
Here, we provide additional details regarding experiments and results in §6. For our prompts, we do not use chat template as it yields worse results in our preliminary experiments.
@@ C.1 Functional Maps
Through the bar plot, we show layer-wise relevance distributions by summing the relevance each head assigns to the question or answer tokens. This reveals whether a particular layer is oriented towards instruction-following or retrieval. In addition, we highlight the top-scoring task and retrieval heads with larger markers.
@@
Figure 6: Functional map of in-context and parametric heads in Mistral-7B-Instruct-v0.3 and Gemma-2-9B-it. Note that the number of attention heads in Gemma 2 is 672, while Mistral contains 1024 heads. The bar plot shows layer-wise sum of heads’ relevance wrt. question tokens or answer tokens located within the context. We highlight the top 40 task and retrieval heads with larger markers.
@@ C.2 Disentangling Functional Roles of In-Context heads
We also compute the sorted in-context scores for Mistral-7B-Instruct-v0.3 and Gemma-2-9B-it. The results are visible in Figure 7. While some task heads show strong in-context behavior, many fall in the moderate range. Notably, Gemma 2 even exhibits parametric task heads, indicating that task heads are not exclusively in-context.
@@ C.3 Task Heads
Table 7 summarizes these results for three different models with roughly similar parameter counts: Llama 3.1 (8B), Mistral v0.3 (7B), and Gemma 2 (9B).
@@
Table 8: Qualitative examples of task FVs patching.
@@ C.4 Parametric Heads
We select random 1,000 examples for each model where their closed-book answer is correct wrt. gold reference (measured by recall with a threshold of 0.7). Then we split them randomly with a proportion of 2.5% train, 2.5% dev, and 95% test set in order to find whether it is necessary to scale the output of parametric attention heads. We consider the scaling factor $\alpha\in[1,1.25,1.5,1.75,2,2.25,2.5,2.75,3]$, maximizing recall on the dev set. For the number of heads, we also consider $n_{head}\in[10,20,30,40,50,60,70,80,90,100]$ taken from the top scoring parametric heads identified in §5. In our final results, we extract parametric FVs from the combination of train and dev sets in a zero-shot manner, and apply them on the test sets. Table 9 shows recall scores on the development set with their optimal scaling factor and number of heads used.
@@
Performance of Parametric FVs
In Table 1, we see that $\mathcal{H}_{\text{param}}$ performs worse compared to other sets. We hypothesize that this is due to how different parametric heads seem to deal with different domains of parametric knowledge, whereas for contextual information this does not seem to be the case. To test this, we patch parametric FVs from cloze-style statements into related question prompts where the attribute is shared. In Table 12, we see that the recall score tend to increase as more heads are used.
Table 12: Zero-shot recall scores for parametric FVs on the test set, separated by attributes. The scaling factor we use is taken from Table 9.
(a) $n_{head}=50$
$\alpha$ Date of birth Place of birth Education Occupation
Parametric FVs
Llama-3.1-8B-Instruct 2 13.33 49.82 51.10 69.36
Mistral-7B-Instruct-v0.3 1.25 24.13 55.68 62.14 51.65
Gemma-2-9B-it 3 09.56 37.20 48.70 66.11
(b) $n_{head}=100$
$\alpha$ Date of birth Place of birth Education Occupation
Parametric FVs
Llama-3.1-8B-Instruct 2 37.41 49.99 52.05 74.59
Mistral-7B-Instruct-v0.3 1.25 35.21 59.15 62.50 72.96
Gemma-2-9B-it 3 11.87 40.62 42.05 63.84
@@ Appendix D Details: Source Tracking
[…]
@@
An optimal decision threshold is then chosen via ROC analysis on a held‑out development subset of $\mathcal{Z}$, selecting the threshold that maximizes the true positive rate while minimizing the false positive rate.
@@
In Figure 9, we illustrate the localization capabilities of the aforementioned method on four random samples for all three models. The aggregated attention maps are plotted as heatmaps, where red colors signify high values and blue colors signify negative colors. Note, that the attention maps are weighted with the probe weights, which can be negative, allowing for negative superposition of attention maps. The beginning of sentence token is receiving some attention weight values due to its usage as attention sinks [84].
Appendix E Additional Results: Resolving Knowledge Conflicts
Table 14 reports the effect of selectively ablating in-context and parametric heads on the NQ-Swap dataset, which contains counterfactuals designed to elicit knowledge conflicts between contextual and parametric information. The ablation is performed incrementally by removing an increasing number of heads identified as belonging to either the in-context or parametric category. When in-context heads are ablated, the model is prevented from relying on contextual information in the input and is forced to depend more heavily on its internal (parametric) knowledge. As shown in the upper part of the table, this generally leads to a substantial increase in recall score wrt. original, non-counterfactual gold answers, indicating that the model gets better at resisting the misleading context and grounds its answers in the stored knowledge. However, the improvement is not strictly monotonic across all models suggesting that the influence of an individual head is not uniform.
Conversely, when parametric heads are ablated (lower part), the model is encouraged to ground its predictions more strongly on the provided context while suppressing the reliance on its internal memory. We again observe substantial gains in recall score wrt. counterfactual open-book answers, although the relationship between the number of the removed heads and the performance is non-monotonic for some models.
Table 14: Ablation of in-context vs parametric heads on NQ-Swap. Recall (%) wrt. gold answer (closed-book) for ablating in-context heads and wrt. counterfactual answer (open-book) for ablating parametric heads.
Ablation (Nr. Heads Removed)
0 20 40 60 80 100
In-Context Heads
Llama-3.1-8B-Instruct 09.5 24.1 24.3 35.1 33.1 16.9
Mistral-7B-Instruct-v0.3 12.0 35.6 37.3 37.6 45.0 45.9
Gemma-2-9B-it 13.3 27.7 23.5 20.3 14.6 09.2
Parametric Heads
Llama-3.1-8B-Instruct 68.7 72.5 73.2 74.2 75.7 76.4
Mistral-7B-Instruct-v0.3 67.6 72.9 75.0 75.3 72.6 62.0
Gemma-2-9B-it 66.6 69.3 71.9 73.5 72.8 73.9
Appendix F Additional Results: Machine Translation
To further demonstrate the generalizability of our approach, we perform evaluation on machine translation using the OPUS Europarl dataset [^12] https://huggingface.co/datasets/Helsinki-NLP/europarl (en-fr, en-es) where we align the first 10,000 examples. The attention heads used for patching are retrieved following the procedure described in §6.1, where we select the top 15 heads exhibiting the highest relevance scores on their key positions for the instruction “Translate into German” in the seed prompt “Translate into German: I love AI research.” We intentionally choose German as neither source nor target language in subsequent evaluations to demonstrate that the extracted feature vector does not encode language-specific information, but rather the general translation instruction.
In the patching setup, we provide the model with inputs of the form “[SENTENCE] $\rightarrow$ and parse the model output following the arrow as the generated translation. The function vector is applied in a zero-shot fashion to induce translation behavior across language pairs. For Llama-3.1-8B-Instruct we use a scaling factor of 1 and for Mistral-7B-Instruct-v0.3 a scaling factor of 2.
[…]
Baseline Random MT Heads Baseline Random MT Heads
Llama-3.1-8B-Instruct 33.2 1.0 31.5 29.5 0.37 24.9
Mistral-7B-Instruct-v0.3 29.2 1.6 18.2 25.2 0.60 18.3
To assess the effectiveness of this approach, we compare it against two baselines (Table 15): (1) direct prompting with “Translate the sentence into [LANGUAGE]: [SENTENCE] $\rightarrow$, and (2) randomly patching attention heads. Our method achieves comparable performance to the explicit prompting baseline demonstrating that the patched task heads reliably induce translation behavior across languages. Notably, this effect is achieved using only 15 of the 1024 attention heads in the Llama-3.1-8B-Instruct and Mistral-7B-Instruct-v0.3 model, underscoring the precision of the identified translation subspace. In contrast, random patching yields near-zero BLEU score, confirming that translation behavior arises from targeted task heads selection.
Unfortunately, the Gemma-2-9B-it model exhibits instability in this setting: it frequently fails to produce outputs adhering to the [SENTENCE] $\rightarrow$ [TRANSLATION] format, making translation parsing unreliable. We leave this issue for future investigation, but we hypothesize that this instability stems from the comparatively small number of attention heads in Gemma 2, which likely leads to stronger superposition effects and less clearly separable functional subspaces compared to models such as Llama 3.1 and Mistral v0.3.
@@ Appendix F Additional Results: Machine Translation
Figure 8: Distributions of logit lens scores for the top 40 retrieval heads in Llama-3.1-8B-Instruct. Shown are the logit lens activations for the ground truth and counterfactual output tokens, comparing cases where the model generates the answer from its parameters (red) versus cases where it retrieves the answer from the context (green) respectively. We observe that retrieval heads exhibit heightened activity when the model relies on context, as indicated by the elevated logit lens scores in green color. While these distributions are 1-dimensional for each head, the probe itself learns a decision boundary in the full 40-dimensional space, where the retrieval signal may be better disentangled. Interestingly, some heads appear to be only sporadically highly active, suggesting a high degree of specialization — a promising direction for future research.
@@
Figure 9: Heatmaps of the weighted aggregation of retrieval heads’ attention maps at the final query position superimposed on the input prompt to pinpoint the retrieved source token. For each model, the aggregated attention maps of the retrieval heads reliably focus on the predicted token in the context, which can be used as cost-effective source tracking. The beginning of sentence token is receiving some attention weight values due to its usage as attention sinks [84].